# Physics-informed Temporal Alignment for Auto-regressive PDE Foundation Models

### v1

@@ B.6 CFDBench-NS
$\frac{\partial}{\partial t}(\rho{\bm{u}})+\nabla\cdot(\rho{\bm{u}^{2}})=-\nabla p+\nabla\cdot\mu\left[\nabla\boldsymbol{u}+(\nabla{\bm{u}})^{\top}\right]$ (18)
@@
B.7 Magnet-Viscous Burgers
@@
Appendix C Training Details
C.1 Hyperparameters
Model DPOT MPP FNO
Batch Size 20 20 24 24 20 20
Gradient Clipping 10000 10000 1 1 1 1
Dropout 0 0 0.1 0.1 0 0
Initial Learning Rate 1e-3 1e-3 1e-3 1e-3 1e-3 1e-3
Optimizer Adam Adam AdamW AdamW Adam Adam
Learning Rate Schedule Cycle Cycle Cycle Cycle Cycle Cycle
Weight Decay 1e-6 1e-6 5e-2 5e-2 1e-6 1e-6
Warmup Epoch 50 50 5 5 50 50
C.2 Model Configuration
@@
Appendix D Supplementary Experimental Results
D.1 Data Efficiency Analysis of Downstream Tasks
@@
D.2 Performance Comparison of PITA and Auto-regressive Methods on Long Trajectory Datasets
@@
D.3 Scaling Experiments
In this section, we analyze the scaling results from Table 1, which are illustrated in Figure 6. It is observed that as the model size increases, the average nRMSE across the five datasets with long trajectories decreases, generally following a scaling law. However, on two larger-scale datasets (FNO-NS-1e-3 and FNO-NS-1e-4) PITA performs less stably on the larger-sized baselines. The main reason may be that the DPOT baseline uses a cyclic learning rate decay incorporating very high learning rates during training. This may lead to gradient explosion or overly aggressive weight updates, affecting the convergence of models with more parameters. Furthermore, DPOT performs finetuning for all model sizes with only 500 epochs and sets the gradient clipping value to 10,000. This prevents PITA from converging stably on larger datasets with larger model sizes. We found that this issue can be mitigated after adjusting the training settings. For example, when we provide sufficient training, PITA shows more stable convergence results, as shown in Table 10 of Appendix D.4. To demonstrate PITA’s plug-and-play performance more fairly, we still follow DPOT’s training settings in the experiments, without making any adjustments to PITA, except for special declarations.
@@ D.3 Scaling Experiments
D.4 Comparative Analysis of Training from Scratch and Finetuning Pretrained Models
@@
D.5 Comparison of PITA and Auto-regressive Methods ($T_{ar}=10$)
@@
D.6 Comparison of Sparse Regression and Other Numerical Algorithms
@@
D.7 Comparison of Computational Cost Between PITA and Autoregressive Methods
PITA differs from conventional autoregressive methods by incorporating a PDE discovery procedure, which introduces additional training-time computation and CPU memory usage. We measured per-batch training time, the extra time incurred by the PDE discovery process (“STRidge time” in Table 13), and CPU memory consumption. Given that GPU memory usage remains invariant relative to the baseline, it is therefore excluded from our measurements. The results are summarized in Table 13 and Figure 7.
@@ D.7 Comparison of Computational Cost Between PITA and Autoregressive Methods
D.8 Influence of the Percentage of Remained Library Terms
@@
D.9 Influence of Accuracy of the Discovered Physics Accuracy on PITA’s Performance
@@
D.10 Performance on Modeling Chaotic Dynamical Systems
@@
D.11 Robustness Analysis of PITA under Noise and Data Masking
@@
Secondly, we evaluate PITA’s capability to reconstruct solutions from incomplete observations. Specifically, 25% of the PDE dataset is randomly masked, and the model is tasked with predicting the full field, including within the occluded regions. As reported in Table 19, PITA significantly outperforms the auto-regressive baseline under these conditions. While all methods experience some performance degradation when confronted with missing data, PITA’s relative drop in accuracy is markedly lower. We attribute this resilience to PITA’s integrated PDE-discovery module, which infers the governing dynamics even from downsampled, partially missing inputs, enabling it to generalize beyond the observed measurements (see Sec. LABEL:Governing_Equations_Discovery). These results demonstrate PITA’s enhanced ability to recover accurate solutions in data-sparse settings, highlighting its potential for deployment in real-world scenarios where measurement gaps are ubiquitous.
@@
D.12 Error Bars of PITA on Long Trajectory Datasets
To ascertain that the improvements of PITA demonstrated in Table 1 stem specifically from its PDE discovery framework and alternating direction optimization algorithm, we evaluate the performance of PITA against autoregressive methods on datasets featuring extended temporal trajectories. Such datasets inherently pose challenges for long-term prediction and downstream tasks due to their temporal complexity. By assessing multiple model scales, we further investigate how variance patterns and reliability metrics correlate with model complexity. As illustrated in Figure 8, the results confirm that PITA’s enhancements are statistically significant and intrinsically tied to its architectural design, rather than stochastic fluctuations.
@@ D.12 Error Bars of PITA on Long Trajectory Datasets
Appendix E Visualizations of Trajectory
@@ E.4 PDEBench-CNS
E.5 Magnet-Viscous Burgers

### v2

@@ C.6 CFDBench-NS
$\frac{\partial}{\partial t}(\rho{\bm{u}})+\nabla\cdot(\rho{\bm{u}^{2}})=-\nabla p+\nabla\cdot\mu\left[\nabla\boldsymbol{u}+(\nabla{\bm{u}})^{\top}\right],$ (18)
@@
C.7 Magnet-Viscous Burgers
@@
Appendix D Training Details
D.1 Model Configuration
@@
D.2 Hyperparameters
Model DPOT MPP FNO
Batch Size 20 20 24 24 20 20
Gradient Clipping 10000 10000 1 1 1 1
Dropout 0 0 0.1 0.1 0 0
Initial Learning Rate 1e-3 1e-3 1e-3 1e-3 1e-3 1e-3
Optimizer Adam Adam AdamW AdamW Adam Adam
Learning Rate Schedule Cycle Cycle Cycle Cycle Cycle Cycle
Weight Decay 1e-6 1e-6 5e-2 5e-2 1e-6 1e-6
Warmup Epoch 50 50 5 5 50 50
Appendix E Supplementary Experimental Results
E.1 Data Efficiency Analysis of Downstream Tasks
@@
E.2 Performance Comparison of PITA and Auto-regressive Methods on Long Trajectory Datasets
@@
E.3 Scaling Experiments
In this section, we analyze the scaling results from Table 1, which are illustrated in Figure 8. It is observed that as the model size increases, the average nRMSE across the five datasets with long trajectories decreases, generally following a scaling law. However, on two larger-scale datasets (FNO-NS-1e-3 and FNO-NS-1e-4) PITA performs less stably on the larger-sized baselines. The main reason may be that the DPOT baseline uses a cyclic learning rate decay incorporating very high learning rates during training. This may lead to gradient explosion or overly aggressive weight updates, affecting the convergence of models with more parameters. Furthermore, DPOT performs finetuning for all model sizes with only 500 epochs and sets the gradient clipping value to 10,000. This prevents PITA from converging stably on larger datasets with larger model sizes. We found that this issue can be mitigated after adjusting the training settings. For example, when we provide sufficient training, PITA shows more stable convergence results, as shown in Table 10 of Appendix E.4. To demonstrate PITA’s plug-and-play performance more fairly, we still follow DPOT’s training settings in the experiments, without making any adjustments to PITA, except for special declarations.
@@ E.3 Scaling Experiments
E.4 Comparative Analysis of Training from Scratch and Finetuning Pretrained Models
@@
E.5 Comparison of PITA and Auto-regressive Methods ($T_{ar}=10$)
@@
E.6 Comparison of Sparse Regression and Other Numerical Algorithms
@@
E.7 Comparison of Computational Cost Between PITA and Auto-regressive Methods
PITA differs from conventional auto-regressive methods by incorporating a PDE discovery procedure, which introduces additional training-time computation and CPU memory usage. We measured per-batch training time, the extra time incurred by the PDE discovery process (“STRidge time” in Table 13), and CPU memory consumption. Given that GPU memory usage remains invariant relative to the baseline, it is therefore excluded from our measurements. The results are summarized in Table 13 and Figure 9.
@@ E.7 Comparison of Computational Cost Between PITA and Auto-regressive Methods
E.8 Influence of the Percentage of Remained Library Terms
@@
E.9 Influence of Accuracy of the Discovered Physics Accuracy on PITA’s Performance
@@
E.10 Performance on Modeling Chaotic Dynamical Systems
@@
E.11 Robustness Analysis of PITA under Noise and Data Masking
@@
Secondly, we evaluate PITA’s capability to reconstruct solutions from incomplete observations. Specifically, 25% of the PDE dataset is randomly masked, and the model is tasked with predicting the full field, including within the occluded regions. As reported in Table 19, PITA significantly outperforms the auto-regressive baseline under these conditions. While all methods experience some performance degradation when confronted with missing data, PITA’s relative drop in accuracy is markedly lower. We attribute this resilience to PITA’s integrated PDE-discovery module, which infers the governing dynamics even from downsampled, partially missing inputs, enabling it to generalize beyond the observed measurements. These results demonstrate PITA’s enhanced ability to recover accurate solutions in data-sparse settings, highlighting its potential for deployment in real-world scenarios where measurement gaps are ubiquitous.
@@
E.12 Error Bars of PITA on Long Trajectory Datasets
To ascertain that the improvements of PITA demonstrated in Table 1 stem specifically from its PDE discovery framework and alternating direction optimization algorithm, we evaluate the performance of PITA against auto-regressive methods on datasets featuring extended temporal trajectories. Such datasets inherently pose challenges for long-term prediction and downstream tasks due to their temporal complexity. By assessing multiple model scales, we further investigate how variance patterns and reliability metrics correlate with model complexity. As illustrated in Figure 10, the results confirm that PITA’s enhancements are statistically significant and intrinsically tied to its architectural design, rather than stochastic fluctuations.
@@ E.12 Error Bars of PITA on Long Trajectory Datasets
Appendix F Visualizations of Trajectory
@@ F.4 PDEBench-CNS
F.5 Magnet-Viscous Burgers